OpenStudy (anonymous):

true or false: there is no solution to the equation csc x = -1 ( I think it's true, but I want to make sure! )

OpenStudy (tkhunny):

Is there a solution to \(sin(x) = -1\)?

OpenStudy (anonymous):

sin x is defined for -1, 0, and 1, so then the answer would be false?

OpenStudy (tkhunny):

Oh, well if sin(x) can give -1, the csc(x) = 1/sin(x) = 1/(-1) = -1 had better exist.